Complete combustion of 0.25moles of an alkane produced 22.4dm³ of CO2 at s.t.p. The molecular formula of the alkane is likely to be???
1
Expert's answer
2020-07-30T07:06:52-0400

Chemical amount of carbon dioxide gas n(CO2) = 22.4/22.4 = 1 mole = n(C atoms)

The general formula of alkane is CnH2n+2. Thus the number of C atoms N(C) = 1/0.25=4. Thus the formula of alkane is C4H10 - butane..